# Distance-based measures and Epsilon-measures for measurement-based quantum resources

### v2

@@ II.1 Quantum Measurements and quantum channels
It should be noted that, usually a factor $1/2$ is added in the diamond distance to normalize it [23]. But we do not use this factor as it is trivial and does not affect any of our results and therefore, is unnecessary. It is known that $\mathcal{D}_{\Diamond}$ satisfies all three conditions (i.e., positivity, symmetric, and triangular inequality) that are required to be satisfied by any distance measure [26].
@@ II.2 Aspects of a resource measure based on set of quantum measurements
A resource measure that satisfies these three conditions is usually said to be a good measure to work with for any convex resource theory. From now on, we assume resource measures to be a continuous function. In our scenario, the objects are the sets of measurements and transformations are the physically realizable maps that transform a set of measurements to another set of measurements in Heisenberg picture.
Given a resource measure $\mathbbm{R}$, one can define a family of resource measures that depend on the parameter $\epsilon$. In the following, we write down the definition of such a measure.
@@ III.1 Transformation of sets of measurements, sets of channels and a distance measure
    where $\mathcal{E}(\rho)\in\mathscr{C}(\mathcal{K},\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ such that $\mathcal{E}(\rho)=\sum_{\lambda}\Phi_{\lambda}(\rho)\otimes\ket{\lambda}\bra{\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{K})$ where {|$\{\ket{\lambda}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{\Lambda}$, for all $j$, $\Psi_{j}\in\mathscr{C}(\mathcal{H}\otimes\mathcal{H}_{\Lambda},\mathcal{H}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ such that $\Psi_{j}(\sigma)=\sum_{i,\lambda}q(i\mid\lambda,j)\ket{i,\lambda}\bra{\lambda}\sigma\ket{\lambda}\bra{i,\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{i}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{I}$, and for all $j$, $\Theta_{j}\in\mathscr{C}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda},\mathcal{H}_{B})$ such that $\Theta_{j}(\omega)=\sum_{y}r(y|i,x,\lambda,j)\ket{y}\bra{x,i,\lambda}\omega\ket{x,i,\lambda}\bra{y}$ | for all $\omega\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{y}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{B}$.

### v3

@@ II.1 Quantum Measurements and quantum channels
It should be noted that, usually a factor $1/2$ is added in the diamond distance to normalize it [23]. But we do not use this factor as it is trivial and does not affect any of our results and therefore, is unnecessary. If $1/2$ factor is included, the upper bound of the diamond distance is $1$ [26]. But as we did not include the $1/2$ factor in the diamond distance, the upper bound is $2$ in our case. It is known that $\mathcal{D}_{\Diamond}$ satisfies all three conditions (i.e., positivity, symmetric, and triangular inequality) that are required to be satisfied by any distance measure [26].
@@ II.2 Aspects of a resource measure based on set of quantum measurements
A resource measure that satisfies these three conditions is usually said to be a good measure to work with for any convex resource theory. From now on, we assume resource measures to be a continuous function. In our scenario, the objects are the sets of measurements and transformations are the physically realizable maps that transform a set of measurements to another set of measurements in Heisenberg picture. Two well-known resource measures that will be used later in this work are the resource robustness and the resource weight, as defined below.
For convex measurement-based resource theories, the resource robustness of an arbitrary set of measurements =$\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
$\begin{split}\mathscr{R}(\mathcal{M})=\min\leavevmode\nobreak\ &\left.r\right.\\ \text{s.t.}\leavevmode\nobreak\ &\left.\mathcal{N}=\{N_{i}=\frac{M_{i}}{1+r}+\frac{r\tilde{M}_{i}}{1+r}\}\in\mathcal{F}_{\mathcal{H}_{A}}\right.\\ &\left.\tilde{\mathcal{M}}=\{\tilde{M}_{i}\}\subset\mathscr{M}(\mathcal{H}_{A}).\right.\end{split}$ (23)
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the set of measurements $\tilde{\mathcal{M}}$.
Similarly, for convex measurement-based resource theories, the resource weight of an arbitrary set of measurements =$\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
$\begin{split}\mathscr{W}(\mathcal{M})=\min\leavevmode\nobreak\ &\left.r\right.\\ \text{s.t.}\leavevmode\nobreak\ &\left.\mathcal{M}=\{M_{i}=\frac{\tilde{N}_{i}}{1+r}+\frac{r\tilde{M}_{i}}{1+r}\}\right.\\ &\left.\tilde{\mathcal{M}}=\{\tilde{M}_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})\right.\\ &\left.\tilde{\mathcal{N}}=\{\tilde{N}_{i}\}\in\mathcal{F}_{\mathcal{H}_{A}}.\right.\end{split}$ (24)
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the set of measurements $\tilde{\mathcal{M}}$.
Given an arbitrary resource measure $\mathbbm{R}$, one can define a family of resource measures that depend on the parameter $\epsilon$. In the following, we write down the definition of such a measure.
@@ III.1 Transformation of sets of measurements, sets of channels and a distance measure
    where $\mathcal{E}(\rho)\in\mathscr{C}(\mathcal{K},\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ such that $\mathcal{E}(\rho)=\sum_{\lambda}\Phi_{\lambda}(\rho)\otimes\ket{\lambda}\bra{\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{K})$ where {|$\{\ket{\lambda}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{\Lambda}$, for all $j$, $\Psi_{j}\in\mathscr{C}(\mathcal{H}\otimes\mathcal{H}_{\Lambda},\mathcal{H}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ such that $\Psi_{j}(\sigma)=\sum_{i,\lambda}q(i\mid\lambda,j)\ket{i,\lambda}\bra{\lambda}\sigma\ket{\lambda}\bra{i,\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{i}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{I}$, and for all $j$, $\Theta_{j}\in\mathscr{C}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda},\mathcal{H}_{B})$ such that $\Theta_{j}(\omega)=\sum_{y,x,i,\lambda}r(y|i,x,\lambda,j)\ket{y}\bra{x,i,\lambda}\omega\ket{x,i,\lambda}\bra{y}$ | for all $\omega\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{y}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{B}$.
@@ III.2 Distance-based measures for measurement-based resources
Proposition 6.
If $\mathbbm{D}=\overline{\mathcal{D}}$, then for any set of measurement $\mathcal{M}\subset\mathscr{M}(\mathcal{H}_{A})$, we have
$\overline{\mathbbm{R}}(\mathcal{M})\leq$ $\min\left\{\frac{2\mathscr{R}(\mathcal{M})}{1+\mathscr{R}(\mathcal{M})},\frac{2\mathscr{W}(\mathcal{M})}{1+\mathscr{W}(\mathcal{M})}.\right\}$ (82)
Proof.
We have $\mathbbm{D}=\overline{\mathcal{D}}$. Suppose in Eq. (23), the minimum occurs for ~==$\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$ and $\mathscr{R}(\mathcal{M})=r^{*}$. Then
$N_{i}=\frac{M_{i}}{1+r^{*}}+\frac{r^{*}M^{*}_{i}}{1+r^{*}}\leavevmode\nobreak\ \forall i$
$or,\leavevmode\nobreak\ N_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{M_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}+\frac{r^{*}\tilde{M}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$ (83)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$ (see Remark 2). Let
$\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (84)
and
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\overline{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (85)
Then observing (from Eq.(23), the assumption [A1] and the assumption [A5]) that $(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$, from Eq. (83), Eq. (84) and Eq. (85), we have for all $\mathcal{H}_{B}$
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\frac{2r^{*}}{1+r^{*}}.$ (86)
Then recalling $\mathscr{R}(\mathcal{M})=r^{*}$, noting the fact that $\mathbbm{D}=\overline{\mathcal{D}}$ and from Eq. (55), we have
$\overline{\mathbbm{R}}(\mathcal{M})$ $\leq\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$
$\leq\frac{2\mathscr{R}(\mathcal{M})}{1+\mathscr{R}(\mathcal{M})}.$ (87)
Suppose in Eq. (24), the minimum occurs for ~==$\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$, $\tilde{\mathcal{N}}=\mathcal{N}^{*}=\{N_{i}^{*}\}$ and $\mathscr{W}(\mathcal{M})=r^{*}$. Then
$M_{i}=\frac{N^{*}_{i}}{1+r}+\frac{r^{*}M^{*}_{i}}{1+r}\leavevmode\nobreak\ \forall i$
$or,\leavevmode\nobreak\ M_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{N^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}+\frac{r^{*}M^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$ (88)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$. Let
$\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (89)
and
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\overline{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (90)
Then (from Eq. (24), the assumption [A1] and the assumption [A5]) that $(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$, from Eq. (88), Eq. (89) and Eq. (90), we have for all $\mathcal{H}_{B}$
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\frac{2r^{*}}{1+r^{*}}.$ (91)
Then recalling $\mathscr{W}(\mathcal{M})=r^{*}$, noting the fact that $\mathbbm{D}=\overline{\mathcal{D}}$ and from Eq. (55), we have
$\overline{\mathbbm{R}}(\mathcal{M})$ $\leq\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$
$\leq\frac{2\mathscr{W}(\mathcal{M})}{1+\mathscr{W}(\mathcal{M})}.$ (92)
Therefore,
$\overline{\mathbbm{R}}(\mathcal{M})\leq$ $\min\left\{\frac{2\mathscr{R}(\mathcal{M})}{1+\mathscr{R}(\mathcal{M})},\frac{2\mathscr{W}(\mathcal{M})}{1+\mathscr{W}(\mathcal{M})}\right\}.$ (93)
Clearly, this constitute a tighter bound of $\overline{\mathbbm{R}}(\mathcal{M})$ than that in Eq. (III.2). However, this may not be the tightest one.